Problem 4 A sample of 15 small bags of the same brand of candies was selected. Assume that the population distribution of bag weights is normal. The weight of each bag was then recorded. The mean weight was 3 ounces with a sample standard deviation0.4 ounces. We would lke to calculate an 80% cofdene interval for the average weight of a sample of size 15 a. (3%) standard error- b. (3%) The critical t value for an 80% confidence interval is crit c (396) EBM- d. (4%) 80% confidence interval for the population average weight of the candies is from An to
Answer #1

Solution :

Given that mean x = 3 , standard deviation s = 0.4 , n = 15

a. standard error = s/sqrt(n)

= 0.4/sqrt(15)

= 0.1033

b. => df = n - 1 = 14

for an 80% confidence interval , the critival t - value is tc = +/- 1.345


c.

=> for an 80% confidence interval , t = 1.345

=> EBM = t*s/sqrt(n)

=> EBM = 1.345*0.4/sqrt(15)

= 0.1389

d.

=> The 80% confidence interval is

=> x - t*s/sqrt(n) to x + t*s/sqrt(n)

=> 3 - 0.1389 to 3 + 0.1389

=> 2.8611 to 3.1389